# Hyperdimensional Probe: Decoding LLM Representations via Vector Symbolic Architectures

### v1

@@ Abstract
Despite their capabilities, Large Language Models (LLMs) remain opaque with limited understanding of their internal representations. Current interpretability methods, such as direct logit attribution (DLA) and sparse autoencoders (SAEs), provide restricted insight due to limitations such as the model’s output vocabulary or unclear feature names. This work introduces Hyperdimensional Probe, a novel paradigm for decoding information from the LLM vector space. It combines ideas from symbolic representations and neural probing to project the model’s residual stream into interpretable concepts via Vector Symbolic Architectures (VSAs). This probe combines the strengths of SAEs and conventional probes while overcoming their key limitations. We validate our decoding paradigm with controlled input–completion tasks, probing the model’s final state before next-token prediction on inputs spanning syntactic pattern recognition, key–value associations, and abstract inference. We further assess it in a question-answering setting, examining the state of the model both before and after text generation. Our experiments show that our probe reliably extracts meaningful concepts across varied LLMs, embedding sizes, and input domains, also helping identify LLM failures. Our work111github.com/Ipazia-AI/hyperprobe advances information decoding in LLM vector space, enabling extracting more informative, interpretable, and structured features from neural representations.
Keywords Information Decoding $\cdot$ LLMs $\cdot$ Interpretability $\cdot$ Probing $\cdot$ Vector Symbolic Architectures
@@ 1 Introduction
Although LLMs excel across tasks, their black-box nature limits interpretability. Recent work has focused on decoding human-interpretable concepts from LLM latent representations (Gurnee and Tegmark, 2023; Park et al., 2023; Zhang et al., 2024). Three main paradigms (Ferrando et al., 2024; Elhage et al., 2021) are currently proposed to inspect model’s residual stream: Supervised Probes, Direct Logit Attribution (DLA), and Sparse Autoencoders (SAEs). Probes are supervised models for task-specific probing objectives that map a model’s vector space to meaningful features (Gurnee and Tegmark, 2023; Marks and Tegmark, 2023; Diego Simon et al., 2024), though their decoding capabilities have been debated (Hewitt and Liang, 2019). DLA projects representations on the LLM’s output vocabulary (Belrose et al., 2023), but this constrains the abstraction to the level of LLM tokens. SAEs learn a sparse proxy representation (Bricken et al., 2023; Lieberum et al., 2024; Kissane et al., 2024), but naming triggered features often suffers from vagueness, verbosity, and data dependence.
@@
Figure 1: We first compress neural representations of the LLM’s next-token task ($F$, blue). Next, we train a neural VSA encoder to map these neural embeddings into a proxy space, VSA encodings structured with input-related concepts ($T$, orange). We then probe the LLM’s embeddings by extracting concepts from VSA encodings using hypervector algebra ($I$, green). This process of concept extraction ultimately enables deeper analysis of the model’s erroneous answers (red).
Vector Symbolic Architectures (VSAs) are a computational framework (Schlegel et al., 2022; Gayler, 1998) inspired by cognitive science (Hawkins, 2021; Piantadosi et al., 2024), increasingly used to map neural representations to human-readable symbols. It has been used for tasks ranging from visual problems such as multi-attribute digit recognition (Frady et al., 2020) and Raven’s progressive matrices (Hersche et al., 2023) to mechanistic interpretability (Knittel et al., 2024).
This work introduces a novel paradigm for decoding information from LLM vector spaces by integrating ideas from symbolic representations and neural probing. We propose Hyperdimensional Probe, a novel approach for decoding human-interpretable information from latent representations of LLMs using VSAs and hypervector operations. We design a supervised and shallow neural network (encoder) to map the LLM’s residual stream into a controlled vector space structured by VSA encodings, projecting its internal activations into human-interpretable and context-relevant concepts. Functioning as a hybrid supervised probe, it harnesses the orthogonality property of VSAs to combine the SAEs’ ability to uncouple superposed subspaces with the interpretability advantages offered by traditional probes. Our method addresses several limitations of prior approaches by: (i) avoiding dependence on the model’s output vocabulary used in DLA, (ii) mitigating the potential confounding effects of task performance in conventional probes, and (iii) eliminating the need for explicit feature naming of SAEs. Section˜4 describes our novel decoding paradigm, and Section˜5 validates it in a controlled setting, also demonstrating its effectiveness in debugging LLM failures. Appendix˜G contrasts our VSA-based results with DLA. Section˜5.3 applies our methodology to the Stanford Question Answering Dataset (Rajpurkar et al., 2016), validating it in a question-answering setting.
@@
Figure˜1 shows our framework, from LLM token mbedding processing, and neural VSA encoder training, to LLM answer explanation. The primary methodological contributions of this work are:
  • Hyperdimensional probe: a novel paradigm for information decoding in LLMs via VSAs;
  • Effective compression of LLM embeddings to probe a wide range of model’s residual stream while reducing the overall computational cost of probing;
  • Enhanced interpretability of neural representations and LLM’s erroneous outputs.
@@ 2 Related work
The latent representations of transformers, also known as residual stream, is a high-dimensional linear vector space that aggregates the outputs of all hidden layers (Elhage et al., 2021).Probing this additive space requires the identification of human-interpretable features across different layers (Ferrando et al., 2024). This investigation is grounded in the linear representation hypothesis, which posits that latent features can be encoded as linear subspaces (Engels et al., 2024), formed and accessed during the forward pass (Park et al., 2023). In recent years, three main paradigms (Ferrando et al., 2024) have been proposed to extract information from this vector space.
@@
is a generic mapping paradigm that maps the model’s residual stream to task-relevant features, measuring how much information about them is embedded (Tenney et al., 2019). Previous works have shown that several features are linearly encoded in transformers, from syntactical information (Hernández López et al., 2023; Diego Simon et al., 2024), to complex concepts such as space-time coordinates (Gurnee and Tegmark, 2023) and truthfulness (Marks and Tegmark, 2023). However, their probing effectiveness is debated due to difficulties in separating information decoding from probe learning (Ferrando et al., 2024; Hewitt and Liang, 2019).
Direct Logit Attribution (DLA)
projects latent representations onto its output vocabulary through the unbundling layer (Geva et al., 2022). This method, also known as Logit Lens (Belrose et al., 2023), interprets outputs as predicted logits at a given point in the forward pass. DLA reveals next-token predictions, assuming that all subsequent layers are bypassed and providing insight into prediction dynamics (Jastrzebski et al., 2017). However, this approach faces key limitations in uncovering features in the LLM vector space, as it relies solely on the next-token representation and is constrained by the model’s token vocabulary. Thus, abstraction is limited, while additional vector transformations might also be necessary (Belrose et al., 2023; Sakarvadia et al., 2023).
@@
use sparse dictionary learning (Olshausen and Field, 1997) to disentangle overlapping subspaces created by superposition (Cunningham et al., 2023). An autoencoder reconstructs the residual stream in an unsupervised fashion, enforcing sparsity in its learned representations. Once trained, these serve as a proxy layer for analysis. SAE activated neurons are interpreted via two strategies: identify representative tokens via DLA (Kissane et al., 2024; Dunefsky et al., 2024); clustering inputs by shared SAE neurons, followed by manual (Jing et al., 2025) or automatic (Bricken et al., 2023; Lieberum et al., 2024) feature naming. While SAEs help addressing superposition, interpreting the resulting features is challenging. Feature naming plays a crucial role in SAE-based analyses but remains problematic: DLA approaches restrict feature abstraction to LLM tokens, whereas example-based approaches can be overly broad and heavily data-dependent.
Our Hyperdimensional Probe functions as a hybrid supervised probe, taking advantage of the orthogonality property of VSAs to combine SAEs’ ability to uncouple superposed subspaces with the interpretability advantages offered by conventional probes. Our controlled vector space mimics the proxy layer of SAEs without requiring a subsequent feature-naming step. Moreover, learning a vector transformation, rather than directly performing a downstream task, as in traditional probes, may better isolate encoded information by reducing task performance confounds. Finally, our method overcomes the key limitation of DLA-based analysis, its dependence on the model’s output vocabulary, by supporting concept sets with unrestricted levels of abstraction, cardinality, and data types. A recent study (Knittel et al., 2024) uses VSAs for the mechanistic interpretability of transformers (GPT-2), showing layer-wise dynamics of neural weights can be seen as VSA-related circuits of word embeddings, attention, and MLP outputs. In contrast, our work decodes the semantics of the model’s residual stream instead of examining the contributions of the model components to its construction during inference.
@@ 3 Background
Vector Symbolic Architectures (VSAs), also known as Hyperdimensional Computing, assume entities or data structures can be represented as random points in a high-dimensional space. Owing to the concentration of measure phenomenon (Ledoux, 2001; Kanerva, 2009), exponentially many distinct concepts can be encoded as nearly orthogonal random vectors. A codebook $\Phi$ maps a predefined set $N$ of concepts to their hypervectors, while orthogonality and simple hypervector operations allow composition into more complex concepts.
@@
We adopt the Multiply-Add-Permute architecture (MAP-Bipolar, MAP-B) from VSAs (Schlegel et al., 2022; Gayler, 1998), using bipolar hypervectors in ${-1,1}^{D}$. Dimensionality $D$, typically $10^{2}$$10^{4}$, depends on the number of concepts (Kanerva, 1988) and representation complexity. MAP-B can theoretically encode $2^{D}$ orthogonal, independent elements (Schlegel et al., 2022). Its codebook $\Phi\in{-1,1}^{n_{c}\times D}$ stores $n_{c}$ atomic concepts as bipolar random vectors, generated deterministically from seeds to ensure orthogonality and independence. Each vector is associated with a concept, and $\Phi$ enables evaluation of representations by comparing them with known vectors. Since MAP-B operates in the bipolar domain, cosine similarity is used (Schlegel et al., 2022).
@@ 4 Hyperdimensional probe [Uncaptioned image]
This section presents our VSA-based paradigm for extracting human-interpretable information from LLM latent representations. In Section˜4.1, we introduce a synthetic corpus of diverse analogies, providing a simple and controlled environment to validate our decoding method. Section˜5.3 applies our methodology in a QA setting, while Appendix˜H discusses other settings. Section˜4.2 then presents the construction of input representations using the hypervector algebra. We then illustrate our three-stage pipeline: (a) processing LLM embeddings (Section˜4.3, $F$ in Figure˜1); (2) the neural VSA encoder that maps embeddings onto our controlled proxy space, yielding VSA encodings (Section˜4.4, $T)$; and (3) the extraction of concepts from this proxy space (Section˜4.5, $I)$.
@@ 4.1 Synthetic corpus
We build a textual dataset to evaluate the key components of our decoding paradigm in a simple, controlled, and interpretable testbed. Using controlled input-completion tasks also allows us to focus LLMs on concepts and their relationships, while testing inputs demanding diverse reasoning from syntactical pattern recognition and key-value association to abstract inference.
@@
This work focuses on analogies, textual inputs representing pairs of concepts connected by the same type of factual, syntactic, or semantic relationship. We collect pairs of analogies from two knowledge bases: Google analogy test set (Mikolov, 2013), and the Bigger Analogy Test Set (BATS, (Gladkova et al., 2016)). These span 44 domains across five distinct categories, covering a wide range of factual and linguistic relationships, including analogies related to factual knowledge (e.g., a country’s currency), semantic relations (e.g., grammatical gender), and morphological modifiers (e.g., verb+men). We also design mathematical analogies using three-digit integers and basic operations such as doubling, cubing, division, and extraction of roots.
@@
where a1 and b1 represent the keys of the two pairs, and a2 and b2 are their corresponding values. For example, Denmark:krone = Mexico:peso for the countries currencies, and queen:king = mother:father for the grammatical gender. Table˜11 and Table˜12 in Appendix˜M show the domains grouped by knowledge base and category, respectively. Some concepts span multiple domains, such as Australia links to Canberra, English, and Australian. These overlaps can help mitigate the confounding effect of memorizing key-value pairs. For our experiments in Section˜5, we further limit confounding effects by using the same pairs but generating a set of textual inputs $(\mathcal{\bar{S}})$ with a verbose template: a1 is to a2 as b1 is to b2 (Figure˜2). Conversely, for training (Section˜4.4), we apply data augmentation strategies on $\mathcal{S}$, such as key-value swapping, effectively tripling the corpus size which results in 395,944 training inputs (Appendix˜M).
@@
4.2 Input representations
This section describes the process of constructing VSA encodings for the training stage. This procedure, illustrated with our textual templates (Equation˜1), generalizes to other templates (e.g., question-answer in Equation˜5) or tasks (e.g., toxicity detection; Appendix˜H) since VSAs and hypervector algebra can encode complex structures across diverse inputs.

### v2

@@ Abstract
Despite their capabilities, Large Language Models (LLMs) remain opaque with limited understanding of their internal representations. Current interpretability methods either focus on input-oriented feature extraction, such as supervised probes and Sparse Autoencoders (SAEs), or on output distribution inspection, such as logit-oriented approaches. A full understanding of LLM vector spaces, however, requires integrating both perspectives, something existing approaches struggle with due to constraints on latent feature definitions. We introduce the Hyperdimensional Probe, a hybrid supervised probe that combines symbolic representations with neural probing. Leveraging Vector Symbolic Architectures (VSAs) and hypervector algebra, it unifies prior methods: the top-down interpretability of supervised probes, SAE’s sparsity-driven proxy space, and output-oriented logit investigation. This allows deeper input-focused feature extraction while supporting output-oriented investigation. Our experiments show that our method consistently extracts meaningful concepts across LLMs, embedding sizes, and setups; uncovering concept-driven patterns in analogy-oriented inference and QA-focused text generation. By supporting joint input–output analysis, this work111github.com/Ipazia-AI/hyperprobe advances semantic understanding of neural representations while unifying the complementary perspectives of prior methods.
Keywords Neural embeddings $\cdot$ Probing $\cdot$ LLMs $\cdot$ Information Decoding $\cdot$ Vector Symbolic Architectures
@@ 1 Introduction
The black-box nature of LLMs restricts the interpretability of their internal representations, motivating efforts to extract human-interpretable features from their vector spaces (Park et al.,, 2023). Such efforts typically fall into two probing perspectives (Ghandeharioun et al.,, 2024): (1) Output Distribution Inspection, and (2) Feature Extraction. These serve different yet complementary goals: feature-extraction methods target latent input features, whereas output-oriented inspection examines misalignment between a model’s internal state and its textual outputs, a task also known as Eliciting Latent Knowledge (ELK). Comprehensive interpretability of LLM vector spaces would however require combining both, revealing high-level input features that may underlie such misalignment.
@@
Figure 1: We first compress model’s internal state using dimensionality-reduction steps on LLM embeddings ($F$, blue). Next, we train a neural VSA encoder to map these compressed neural embeddings into a bounded proxy space: VSA encodings ($T$, orange). We then perform concept extraction by querying this proxy space using hypervector algebra ($I$, green). Lastly, extracted concepts ultimately enables deeper analysis on model’s internal state and its outputs (red).
Probing approaches for output distribution inspection, such as lens-style (Nostalgebraist,, 2020; Belrose et al.,, 2023) and patching-based methods (Ghandeharioun et al.,, 2024), project model’s embeddings space into its output vocabulary space, limiting the inspection to layer-wise projected next-token predictions in ELK-centered, output-constrained settings. In contrast, input-oriented feature extraction is currently performed using Sparse Autoencoders (SAEs), and classification-oriented supervised probes. The supervised probes map the embeddings space into latent features to assess how much information about them is encoded using a top-down probing strategy (Gurnee and Tegmark,, 2023; Marks and Tegmark,, 2023; Diego Simon et al.,, 2024). This mapping paradigm faces two key challenges, as its probes are: tailored to task-specific features, limiting generalizability; and trained directly on the final features, making it challenging to distinguish information decoding from probe learning. On the other hand, SAEs provide general-purpose feature extraction via a bottom-up probing strategy on a layer-localized proxy space. This, however, reveals an unbounded feature space that requires naming and filtering for interpretability. Meanwhile, comprehensive interpretability of neural embeddings would require integrating input-driven feature extraction with output-oriented analysis to identify the internally represented input–output features during LLM inference.
Our work uses input-oriented feature extraction for output-oriented analysis, examining which concepts are encoded in LLM vector space across two settings: (1) completing analogy-style inputs, and (2) QA-oriented text generation. We address key questions about LLM reasoning: whether models represent all in-context concepts, key–value associations, or only the output concept (RQ1–RQ3), and whether concept-oriented patterns emerge across input types (RQ4) and across diverse LLMs (RQ5; Section˜5.2). During text generation, we further investigate latent patterns by uncovering how many question- and answer-related concepts appear before and after generation (RQ6; Section˜6).
Unfortunately, existing neural-embedding interpretation methods restrict this two-sided investigation, due to their constraints on how latent features are defined. Our work introduces Hyperdimensional Probe, a novel probing paradigm that combines ideas from symbolic representations and neural probing. Inspired by the sparsity constraints of Sparse Autoencoder (SAEs), we exploits properties of Vector Symbolic Architectures (VSAs) and hypervector algebra to overcome major constraints in prior analysis. Functioning as a hybrid supervised probe, our method combines the bounded, interpretable feature space of standard supervised probes with the queryable proxy space of SAE-based methods. It also removes the single-token representation constraint of conventional output-oriented inspection, enabling feature sets with unrestricted abstraction and symbol sources.
@@
Figure˜1 shows our framework, from embedding processing, and neural VSA encoder training, to output-oriented inspection. Our work presents both methodological and experimental contributions:
  • Vector Symbolic Architectures for neural probing: latent features of LLM vector space can be expressed through VSA encodings and hypervector algebra across diverse LLMs;
  • Hyperdimensional probe: a novel probing paradigm that performs input-aligned feature extraction while supporting output-oriented interpretability (Section˜4);
  • Effective compression of LLM embeddings: enables probing the full residual stream (Section˜4.3), removing the need for the layer-selection stage common in earlier work;
  • Concept-oriented insights into LLM inference: Section˜5.2 reveals differences in conceptual richness across inputs and models, while Section˜6 exposes patterns during text generation.
@@ 2 Related work
The latent representations of transformers, also known as residual stream, is a high-dimensional linear vector space that aggregates the outputs of all hidden layers (Elhage et al.,, 2021). In recent years, three main approaches have been used to study the features encoded in this vector space (Ferrando et al.,, 2024): (1) supervised probes, (2) SAEs for input-focused feature extraction, and (3) logit-based methods for output distribution inspection.
@@
is a generic mapping paradigm that maps neural embeddings to task-relevant input features, measuring how much information about them is embedded (Tenney et al.,, 2019). Conventional probes are tailored and trained directly on task-specific features, from syntactical information (Hernández López et al.,, 2023; Diego Simon et al.,, 2024), to space-time coordinates (Gurnee and Tegmark,, 2023) and truthfulness (Marks and Tegmark,, 2023). However, tailoring probes to specific tasks limits generalizability, whereas directly learning targeted features complicates distinguishing actual information decoding from probe-induced learning (Hewitt and Liang,, 2019).
@@
provide input-oriented feature extraction using a proxy space learned via sparse dictionary learning (Olshausen and Field,, 1997), uncovering superposed latent features (Cunningham et al.,, 2023). An autoencoder reconstructs the residual stream in an unsupervised fashion, enforcing sparsity in its learned representations. Once trained, these serve as a proxy layer for analysis. SAE activated neurons are interpreted via two strategies: identify representative tokens via logit-based methods (Kissane et al.,, 2024; Dunefsky et al.,, 2024); clustering inputs by shared SAE neurons, followed by manual (Jing et al.,, 2025) or automatic (Bricken et al.,, 2023; Lieberum et al.,, 2024) feature naming. The shortcomings of SAEs that hinder output-oriented investigation are threefold: (1) their unbounded feature space yields set of latent features that are difficult to control and align; (2) their feature-naming process restricts semantic grounding to either output-vocabulary labels or ambiguous, data-dependent descriptors; and (3) their dependence on layer-localized representations requires identifying a task-specific optimal hidden layer.
Logit Attribution
focuses on output distribution inspection by projecting model’s embedding space onto its output vocabulary through either (1) the model’s unembedding matrix, also known as Direct Logit Attribution (DLA; Logit Lens by Nostalgebraist,, 2020); (2) learned affine transformation (Tuned Lens by Belrose et al.,2023); or (3) patched LLM inference (Patchscope by Ghandeharioun et al.,2024). This logit-based paradigm offers insight into the output distribution (Jastrzebski et al.,, 2017) by generating projected logits at a chosen point in the forward pass, revealing next-token predictions under the assumption that all subsequent layers are bypassed. However, logit-attribution approaches hinder input-aligned feature extraction by (1) relying sorely on token-level surface features, which constrains probing to ELK-style, output-aligned analysis, and (2) restricting features to single-token representations, limiting semantic-oriented analysis of higher-level abstractions.
Functioning as a hybrid supervised probe, Hyperdimensional Probe exploits VSAs and hypervector algebra to unify the diverse perspectives of prior methods. Our novel paradigm integrates (1) the top-down interpretability of conventional probes, (2) the SAE’s ability to learn a sparsity-driven proxy space, and (3) a higher-level, jointly input–output analytic perspective that goes beyond conventional logit-based methods. By leveraging VSA-based representations (Section˜3), we avoid the unbounded feature spaces of SAEs and the need for post-hoc feature naming and filtering. In addition, our embedding-ingestion algorithm (Section˜4.3) removes the requirement to identify an optimal hidden layer of prior methods. Our approach also mitigates the dichotomy between information decoding and probe-induced learning by learning a transformation from LLM embeddings to a controlled proxy space (Section˜4.4) rather than predicting latent features directly. Lastly, it addresses the core limitations of conventional output distribution investigations by supporting feature sets of arbitrary abstraction, cardinality, and symbolic origin. Thus, our novel methodology combines the strengths of prior approaches for joint input-output feature extraction with deeper output-oriented analysis by:
  1. 1.
    defining a compositional bounded feature space (probes) rather than unbounded (SAEs);
  2. 2.
    learning a sparse proxy space (SAEs) rather than targeting task-specific feature (probes);
  3. 3.
    querying a proxy space holistically (SAEs), in contrast to classification-based probes;
  4. 4.
    inspecting the LLM vector space without introducing layer selection (probes, logit, SAEs);
  5. 5.
    targeting concept-oriented latent features (VSAs) rather than token-aligned logit features.
@@ 3 Background
Vector Symbolic Architectures (VSAs; Hyperdimensional Computing) represent entities as random points in a high-dimensional space. Leveraging the concentration of measure phenomenon (Ledoux,, 2001; Kanerva,, 2009), exponentially many distinct concepts can be represented as nearly orthogonal random vectors. A codebook $\Phi$ maps a predefined set of concepts to their hypervector, while orthogonality and hypervector operations allow composition into more complex representations.
@@
We adopt the Multiply-Add-Permute architecture (MAP-Bipolar, MAP-B) from VSAs (Schlegel et al.,, 2022; Gayler,, 1998), using bipolar hypervectors in ${-1,1}^{D}$. Dimensionality $D$, typically $10^{2}$$10^{4}$, depends on the number of concepts (Kanerva,, 1988) and representation complexity. MAP-B can theoretically encode $2^{D}$ orthogonal, independent elements (Schlegel et al.,, 2022). Its codebook $\Phi\in{-1,1}^{n_{c}\times D}$ stores $n_{c}$ atomic concepts as bipolar random vectors, generated deterministically from seeds to ensure orthogonality and independence. Each vector is associated with a concept, and $\Phi$ enables evaluation of representations by comparing them with known vectors. Since MAP-B operates in the bipolar domain, cosine similarity is used (Schlegel et al.,, 2022).
@@ 4 Hyperdimensional probe [Uncaptioned image]
This section introduces our VSA-based framework for extracting information from the neural embeddings of LLMs on analogy-style completion tasks (RQ1–RQ5). Section˜4.1 presents an analogy-style corpus as a controlled testbed for testing our method on inputs requiring varied reasoning. Section˜4.2 explains how training examples are built using hypervector algebra. We next outline our three-stage pipeline: (1) processing embeddings (Section˜4.3, $F$ in Figure˜1); (2) mapping them into a controlled proxy space via a neural VSA encoder, producing VSA encodings (Section˜4.4, $T)$; and (3) extracting concepts from these encodings (Section˜4.5, $I)$. Finally, Section˜6 applies the methodology to text generation (RQ6), and Appendix˜H discusses its broader applicability.
@@ 4.1 Synthetic corpus
We build a textual dataset to evaluate the core components of our approach within a controlled, and interpretable testbed. Using analogy-style tasks guides LLMs toward concepts and their relations, testing diverse reasoning from syntactic patterns and key-value associations to abstract inference.
@@
This work focuses on analogies, textual inputs representing pairs of concepts connected by the same type of factual, syntactic, or semantic relationship. We collect pairs of analogies from two knowledge bases: Google analogy test set (Mikolov,, 2013), and the Bigger Analogy Test Set (BATS, (Gladkova et al.,, 2016)). These span 44 domains across five distinct categories, covering a wide range of factual and linguistic relationships, including analogies related to factual knowledge (e.g., a country’s currency), semantic relations (e.g., grammatical gender), and morphological modifiers (e.g., verb+men). We also design mathematical analogies using three-digit integers and basic operations such as doubling, cubing, division, and extraction of roots.
@@
where a1 and b1 represent the keys of the two pairs, and a2 and b2 are their corresponding values. For example, Denmark:krone = Mexico:peso for the countries currencies, and queen:king = mother:father for the grammatical gender. Table˜11 and Table˜12 in Appendix˜M show the domains grouped by knowledge base and category, respectively. Some concepts span multiple domains, such as Australia links to Canberra, English, and Australian. These overlaps can help mitigate the confounding effect of memorizing key-value pairs. For our experiments in Section˜5, we further limit confounding effects by using the same pairs but generating a set of textual inputs $(\mathcal{\bar{S}})$ with a verbose template: a1 is to a2 as b1 is to b2. Conversely, for training (Section˜4.4), we apply data augmentation strategies on $\mathcal{S}$, such as key-value swapping, effectively tripling the corpus size which results in 395,944 training inputs (Appendix˜M).
@@
4.2 Training examples
This section describes the process of building VSA-based representations for training input. This procedure, illustrated with our textual templates (Equation˜1), generalizes to other templates (e.g., question-answer) or tasks (e.g., toxicity detection; Appendix˜H) since VSAs and hypervector algebra can encode complex structures across diverse inputs.